# LLM one-shot style transfer for Authorship Attribution and Verification

### v2

@@ Abstract
Computational stylometry studies writing style through quantitative textual patterns, enabling applications such as authorship attribution, identity linking, and plagiarism detection. Existing supervised and contrastive approaches often rely on datasets with spurious correlations, conflating style with topic. Despite the relevance of language modeling to these tasks, the pre-training of modern large language models (LLMs) has been underutilized in general authorship analysis. We introduce an unsupervised LLM-based framework that leverages CLM pre-training and in-context learning to measure style transferability between texts using log-probabilities. This approach avoids explicit supervision and topic bias while exploiting the rich stylistic knowledge encoded in LLMs. Our method substantially outperforms unsupervised prompting-based baselines of similar scale and exceeds contrastively trained models when controlling for topical overlap. Performance scales consistently with model size and, in the case of authorship verification, with an additional mechanism that increases test-time computation; enabling flexible trade-offs between computational cost and accuracy.
@@ 1 Introduction
Computational stylometry is the quantitative and computational analysis of writing style, utilizing patterns in vocabulary, syntax, punctuation, and character sequences to identify, verify, or cluster authors based on textual data [daelemans2013ExplanationComputational]. Authorship analysis has been widely applied across diverse domains, with applications spanning both forensic and humanities contexts. In forensics, stylistic and linguistic patterns have been leveraged to analyze terrorist discourse [abbasi2005ApplyingAuthorship], link user identities across online platforms [shu2017UserIdentity], detect plagiarism [stamatatos2011Plagiarismauthorship], and identify coordinated disinformation campaigns [stiff2022Detectingcomputergenerated]. Beyond these security-driven settings, authorship methods also play a key role in the humanities, supporting tasks such as the attribution of pseudepigraphic texts [koppel2013Automaticallyidentifying] and the analysis of anonymous literary works [stover2016Computationalauthorship].
Authorship analysis is a long-standing challenge in natural language processing, with early approaches that take advantage of hand-crafted stylometric features [stamatatos2016AuthorshipVerification, lagutina2019SurveyStylometric]. The advent of pre-trained transformers [vaswani2017AttentionAll, devlin2019BERTPretraining] led to approaches using pre-trained encoder-only transformers [fabien-etal-2020-bertaa, zangerle2023Overviewmultiauthor], allowing more sophisticated encoding of semantic information, yet these models often struggle in the absence of topical correlations [zangerle2023Overviewmultiauthor].
A different supervised approach in the literature involves fine-tuning individual language models for candidate authors [huang2025Attributingauthorship, barlas2020CrossDomainAuthorship], measuring the probability assigned to texts by the language model. While this approach tries to leverage the pre-training of LLMs, it is expensive to apply, and not useful in low resource contexts like authorship verification tasks.
@@
We propose a novel unsupervised approach to authorship attribution and verification based on the extensive CLM pre-training of modern LLMs and their in-context learning capabilities. These methods are explained in fig.˜1 and section˜3, and rely on the log-probabilities of an LLM to measure style transferability from one text to another (OSST score).
Roughly speaking, we consider the task of styling a version in neutral style of a text (obtained via LLM prompting) to recover the original one. This task can be performed by an LLM, and, further, we can provide it with in-context few-shot examples. Given a single in-context one-shot example, the average LLM log-probabilities of the target text (OSST score) reflect how helpful the style of the one-shot example was. Therefore, we use this metric to separate between texts written by the same and different authors. Our approach is completely zero-shot in the case of closed-set attribution and only requires the selection of a decision boundary for open-set attribution and verification.
@@
    We propose a novel unsupervised approach to authorship attribution and verification based on the extensive CLM pre-training of modern LLMs and their few-shot in-context learning capabilities.
@@
    We empirically validated their performance for authorship attribution and verification across several datasets and compared with a contrastive learning and two LLM prompting baselines.
@@ 2 Background
In this section, we provide the necessary background for the article and review the most important and recent work in the field.
@@ 2.1 Authorship tasks
Given a set of texts $T_{a}$ known to be written by the author $a$, the task is to make a binary decision on whether a target text $t$ was written by $a$. A common setting is to have a single text in $T_{a}$, making this a symmetric task of deciding whether two texts are written by the same author.
@@ 2.2 Available data
PAN competitions have similarly contributed standardized AV datasets. The 2020 and 2021 editions [bischoff2020ImportanceSuppressing, bevendorff2020PAN20Authorship] centered on fanfiction, with the latter introducing an out-of-distribution test set featuring unseen authors and novel topics. In 2022 [stamatatos2022PAN22Authorship], the AV task expanded to encompass diverse discourse types, including essays, emails, text messages, and business memos—highlighting the impact of medium on writing style and increasing task difficulty.
Additionally, we leverage AV datasets derived from the PAN style-change detection tasks. These require models to determine whether consecutive paragraphs are authored by the same individual. We consider data from the 2022 [zangerle2022PAN22Authorship], 2023 [zangerle2023PAN23MultiAuthor], and 2024 [zangerle2024PAN24MultiAuthor] competitions. The 2022 dataset features StackExchange posts, while the 2023 and 2024 datasets focus on Reddit, specifically curating subsets where input texts revolve around the same topic. This setting limits reliance on semantic cues and accentuates the need for nuanced stylistic detection.
@@ 2.3 Authorship methods
More recently, there has been growing interest in learning author embeddings, where each text is mapped into a vector representation capturing stylistic properties. These methods often rely on contrastive learning [ai2022WhodunitLearning, huertas-tato2024PARTPretrained, huertas-tato2024Understandingwriting] and assess authorship via similarity in the embedding space. However, such approaches remain dependent on labeled data that is often confounded with topical similarity.
@@ 2.4 Causal Language Modeling and few-shot learning
Causal Language Modeling (CLM) refers to conditioning on prior tokens to predict the next token in a unidirectional, autoregressive manner, ie, modeling $P(x_{t}\mid x_{<t})$ [bengio2003neuralprobabilistic]. It is the pre-training objective used by GPT-style decoder-only models [radford2019language]. A landmark discovery in large language models (LLMs) is their few‑shot in‑context learning ability [wei2022Emergentabilities]. GPT 3 (175B parameters) was shown to perform new tasks with just a few textual examples provided in the prompt, without any gradient updates or fine-tuning, demonstrating competitive performance with fully fine-tuned systems [brown2020LanguageModels]. This emergent capability highlights that massive autoregressive pre-training endows models with flexible, task‑agnostic generalization using only prompt examples.
In the context of this work, we hypothesize that models trained using the CLM objective inherently learn about style, as it is a key component in generating coherent continuations of a text. Furthermore, our proposed method relies on LLMs having the ability to transfer the style of one text to another through one-shot demonstrations.
@@ 3 One-shot Style Transfer (OSST) with LLMs for authorship attribution and verification
In this section we explain in-depth the main components and intuitions behind our OSST-based methods for AA and AV. A high-level overview of all the components is given in fig.˜1.
@@
Our methods require a way to generate neutrally styled versions of a given text. To do this, we simply prompt an instruction-tuned LLM to rewrite the input text in a neutral style, using the prompt shown in LABEL:lst:neutralize-prompt. More formally, given a text $t$, we consider the random variable $\mathrm{N}_{M,\theta}(t)$ of generating a neutral text with LLM $M$ and sampling parameters $\theta$. For the rest of the work, we omit the subscript and abuse the notation to signify a sample of the random variable.
@@
where the prediction is the candidate with the highest score.
@@
    Zero-shot baseline (nOSST-ZS). We normalize the $\mathrm{OSST}$ score by comparing it with the log-probabilities of the zero-shot problem, without any in-context example. This is shown graphically in fig.˜1, and mathematically in the following formula:
@@
    This normalization step presents some problems. For example, fig.˜3 shows the median $\mathrm{nOSST}_{zs}$ score by token position in the PAN23 style change dataset. As we can see, the first few tokens have a much higher median value. This seems natural, as in the zero-shot problem the task is not completely clear during the first tokens. This presents a challenge in comparing texts of different lengths. We also observe differences between models and datasets.
@@ 4.1 Authorship attribution
We evaluated our models in established authorship attribution benchmarks in the fiction domain, including PAN12 [juola_2012_3713273], PAN18 [kestemont_2018_3737849], and PAN19 [kestemont_2019_3530313], and in the corporate email domain with PAN11 [argamon_2011_3713246]. Our main baseline is STAR [huertas-tato2024Understandingwriting], the strongest contrastive learning method reported in prior work. Additionally, we do not compare against LLM prompting approaches such as LIP [huang2024CanLarge], as these methods require large context windows in AA, resulting in markedly lower performance at moderate model sizes.
@@
Table˜1 summarizes results in the open-set setting, where all methods operate in a zero-shot regime and produce similarity scores between text pairs. Scores are standardized across candidate texts to ensure comparability between target texts. A small held-out subset is used to select the decision threshold. Our models exhibit robust performance in the open-set scenario, with all parameter sizes surpassing STAR. Notably, OSST-based models demonstrate less performance degradation relative to closed-set attribution, as measured by accuracy.
Table 1: Open-set authorship attribution evaluation. Best results are highlighted in bold. BinF1Macro denotes the macro-averaged F1 score for the binary classification of true author presence among candidates, while the average rank of the correct candidate is calculated only for instances where the true author is included. All similarity scores are standardized across candidate texts
@@ 4.2 Authorship verification
As in authorship attribution (AA), we compare against the STAR baseline [huertas-tato2024Understandingwriting]. Given that AV requires less context in LLM prompts than AA, we additionally include LIP [huang2024CanLarge] and PromptAV [hung2023WhoWrote] with a 24B Mistral LLM as zero-shot baselines. For STAR and OSST-based methods, we select a decision boundary with 2000 examples from the training set.
@@
The results on authorship verification datasets are summarized in figs.˜4 and 5. Several key patterns emerge: (i) Performance on challenging tasks:nOSST consistently surpasses STAR on settings that control for topical correlations (the third tasks in PAN23ST and PAN24ST), even in its zero-shot (ZS) configuration with an 8B model. In contrast, STAR maintains an advantage in datasets with strong topic biases. (ii) Comparison to LLM prompting:At comparable model scales (up to 24B parameters), nOSST methods substantially outperform LLM-based prompting approaches. (iii) Impact of the RS variant:The RS variant of nOSST generally improves performance across multiple datasets, though a few exceptions remain. (iv) Scaling behavior:Performance exhibits a clear scaling trend: larger nOSST base models tend to achieve higher accuracy, with only minor deviations from monotonic improvement.
@@ 4.5 Multi-language performance in authorship attribution
In fig.˜8, we present the average accuracy and rank of correct candidate of the proposed OSST-based method across various languages and backbone models for authorship attribution (AA), using the PAN18 and PAN19 multi-language AA datasets. STAR is included as a baseline for comparison. As anticipated, performance declines for all models outside of English, with the OSST-based method consistently outperforming STAR by a larger margin in non-English settings. This disparity can be attributed to the limited multilingual training of STAR, whereas OSST benefits from pre-training on large-scale cross-lingual corpora. These results highlight OSST’s superior generalization and robust multilingual capabilities in authorship attribution tasks.
@@ 5 Conclusions
We introduced OSST-based methods, a family of unsupervised approaches to authorship attribution and verification that leverage the extensive causal language modeling pre-training of modern LLMs. This paradigm enables scalable authorship analysis without relying on labeled data prone to topical biases.
@@ Appendix F Do we need neutral examples?
Figure 10: Comparison between using (nOSST) and not using (nOSST-NN) a neutral version to compute log-probs

### v3

@@ Abstract
Computational stylometry studies writing style through quantitative textual patterns, enabling applications such as authorship attribution, identity linking, and plagiarism detection. Existing supervised and contrastive approaches often rely on datasets with spurious correlations, conflating style with topic. Despite the relevance of language modeling to these tasks, the pre-training of modern large language models (LLMs) has been underutilized in general authorship analysis. We introduce an unsupervised framework that uses the log-probabilities of an LLM to measure style transferability between two texts. This framework takes advantage of the extensive CLM pre-training and in-context capabilities of modern LLMs. Our approach avoids explicit supervision with spuriously correlated data. Our method substantially outperforms unsupervised prompting-based baselines at similar model sizes and exceeds contrastively trained models when controlling for topical overlap. Our framework’s performance improves with model size. In the case of authorship verification, we present an additional mechanism that increases test-time computation to improve accuracy; enabling flexible trade-offs between computational cost and task performance.
@@ 1 Introduction
Computational stylometry is the quantitative and computational analysis of writing style, using patterns in vocabulary, syntax, punctuation, and character sequences to identify, verify, or cluster authors based on textual data [daelemans2013ExplanationComputational]. Authorship analysis has been widely applied across diverse domains, with applications spanning both forensic and humanities contexts. In forensics, stylistic and linguistic patterns have been used to analyze terrorist discourse [abbasi2005ApplyingAuthorship], link user identities across online platforms [shu2017UserIdentity], detect plagiarism [stamatatos2011Plagiarismauthorship], and identify coordinated disinformation campaigns [stiff2022Detectingcomputergenerated]. Authorship methods also play a key role in the humanities, supporting tasks such as the attribution of pseudepigraphic texts [koppel2013Automaticallyidentifying] and the analysis of anonymous literary works [stover2016Computationalauthorship].
Authorship analysis is a long-standing challenge in natural language processing, with early approaches that take advantage of hand-crafted stylometric features [stamatatos2016AuthorshipVerification, lagutina2019SurveyStylometric]. The advent of pre-trained transformers [vaswani2017AttentionAll, devlin2019BERTPretraining] led to the use encoder-only models [fabien-etal-2020-bertaa, zangerle2023Overviewmultiauthor]. Yet, such fine-tuned models often struggle in the absence of topical correlations [zangerle2023Overviewmultiauthor].
A different supervised approach in the literature involves fine-tuning individual language models for candidate authors [huang2025Attributingauthorship, barlas2020CrossDomainAuthorship], measuring the probability assigned to texts by the language model. While this approach tries to take advantage of the pre-training of LLMs, it is expensive to apply, and not useful in low resource contexts like authorship verification tasks.
@@
We propose a novel unsupervised approach to authorship attribution and verification that leverage the extensive CLM pre-training of modern LLMs and their in-context learning capabilities. For a given text $t$, we obtain a version $N(t)$ in neutral style via LLM prompting. Next, we consider the task of predicting the styled text $t$ from its neutral version $N(t)$. This task can be performed by an LLM, and, further, we can provide it with in-context few-shot examples. Given a single in-context one-shot example, the average of the LLM log-probabilities of the target text $t$ (or OSST score) reflects how helpful the style of the one-shot example was. Therefore, we use this metric to measure style similarity between the texts. The OSST score supports zero-shot closed-set attribution, though it requires the selection of a decision boundary for open-set attribution and verification. We provide a more in-depth explanation in fig.˜1 and section˜3.
@@
    We propose a novel unsupervised approach to authorship attribution and verification taking advantage of the extensive CLM pre-training of modern LLMs and their in-context learning capabilities.
@@
    We empirically validated its performance for authorship attribution and verification across several datasets and compared with a contrastively trained and two prompting-based baselines.
@@ 2.1 Authorship tasks
Given a set of texts $T_{a}$ known to be written by the author $a$, the task is to make a binary decision on whether a target text $t$ was written by $a$. A common setting is to have a single text in $T_{a}$, making this the symmetric task of deciding whether two texts are written by the same author.
@@ 2.2 Available data
PAN competitions have similarly contributed standardized AV datasets. The 2020 and 2021 editions [bischoff2020ImportanceSuppressing, bevendorff2020PAN20Authorship] centered on fanfiction, with the latter introducing an out-of-distribution test set featuring unseen authors and novel topics. In 2022 [stamatatos2022PAN22Authorship], the AV task expanded to encompass diverse discourse types, including essays, emails, text messages, and business memos.
Additionally, we leverage AV datasets constructed from style-change detection tasks. These require models to determine whether consecutive paragraphs are authored by the same person. We consider data from the 2022 [zangerle2022PAN22Authorship], 2023 [zangerle2023PAN23MultiAuthor], and 2024 [zangerle2024PAN24MultiAuthor] competitions. The 2022 dataset features StackExchange posts, while the 2023 and 2024 datasets focus on Reddit, specifically curating subsets where input texts revolve around the same topic. This setting limits reliance on semantic cues and accentuates the need for nuanced stylistic detection.
@@ 2.3 Authorship methods
More recently, there has been growing interest in learning author embeddings, where each text is mapped into a vector representation capturing stylistic properties. These methods often rely on contrastive learning [ai2022WhodunitLearning, huertas-tato2024PARTPretrained, huertas-tato2024Understandingwriting] and assess authorship via similarity in the embedding space. However, such approaches remain dependent on labeled data, often strongly correlated with topical similarity.
@@ 2.4 Causal Language Modeling and few-shot learning
Causal Language Modeling (CLM) refers to predicting the next token based on prior tokens in a unidirectional, autoregressive manner, ie, modeling $P(x_{t}\mid x_{<t})$ [bengio2003neuralprobabilistic]. It is the pre-training objective used by GPT-style decoder-only models [radford2019language]. A landmark discovery in large language models (LLMs) is their few‑shot learning ability [wei2022Emergentabilities]. GPT 3 (175B parameters) was shown to perform new tasks with just a few textual examples provided in the prompt, without any explicit gradient updates or fine-tuning, demonstrating competitive performance with fully fine-tuned systems [brown2020LanguageModels]. This emergent capability highlights that massive autoregressive pre-training endows models with flexible, task‑agnostic generalization using only prompt examples.
We hypothesize that models trained using the CLM objective learn about style, as it is a key component in generating coherent continuations of a text. Furthermore, our proposed method relies on LLMs having the ability to transfer the style of one text to another through one-shot demonstrations.
@@ 3 One-shot Style Transfer (OSST) with LLMs for authorship attribution and verification
In this section we provide an in-depth explanation of the main components and intuitions behind our OSST-based methods for AA and AV. A high-level overview of all the components is given in fig.˜1.
@@
Our methods require a way to generate neutrally styled versions of a given text. To do this, we prompt an instruction-tuned LLM to rewrite the input text in a neutral style, using the template shown in LABEL:lst:neutralize-prompt. More formally, given a text $t$, we consider the random variable $\mathrm{N}_{M,\theta}(t)$ of generating a neutral text with LLM $M$ and sampling parameters $\theta$. For the rest of the work, we omit the subscript and abuse the notation to signify a sample of the random variable.
@@
and the prediction is the candidate with the highest score.
@@
    Zero-shot baseline (nOSST-ZS). We normalize the $\mathrm{OSST}$ score by comparing it with the log-probabilities of the zero-shot problem, without any in-context example. This is shown graphically in fig.˜1, and mathematically as follows:
@@
    This normalization step presents some problems. Figure˜3 shows the median $\mathrm{nOSST}_{zs}$ score by token position in the PAN23 style change dataset. As we can see, the first few tokens have a much higher median value. This seems natural, as in the zero-shot problem the task is not completely clear during the first tokens. This presents a challenge in comparing texts of different lengths. We also observe differences between models and datasets.
@@ 4.1 Authorship attribution
We evaluated our models in established authorship attribution benchmarks in the fiction domain, including PAN12 [juola_2012_3713273], PAN18 [kestemont_2018_3737849], and PAN19 [kestemont_2019_3530313], and in the corporate email domain with PAN11 [argamon_2011_3713246]. Our main baseline is STAR [huertas-tato2024Understandingwriting], the strongest contrastive learning method reported in prior work. We do not compare against LLM prompting approaches such as LIP [huang2024CanLarge], as these methods require large context windows in AA, resulting in markedly lower performance at moderate model sizes.
@@
Table˜1 summarizes results in the open-set setting, where all methods operate in a zero-shot regime and produce similarity scores between text pairs. Scores are standardized across candidate texts (for all models) to improve comparability of scores for different target texts. A small held-out subset is used to select the decision threshold. Our models exhibit robust performance in the open-set scenario, with all parameter sizes surpassing STAR. Notably, OSST-based models demonstrate less performance degradation relative to closed-set attribution, as measured by accuracy.
Table 1: Open-set authorship attribution evaluation. Best results are highlighted in bold. BinF1Macro denotes the macro-averaged F1 score for the binary classification of true author presence among candidates, while the average rank of the correct candidate is calculated only for instances where the true author is included. All similarity scores are standardized across candidate texts.
@@ 4.2 Authorship verification
As in authorship attribution (AA), we compare against the STAR baseline [huertas-tato2024Understandingwriting]. We include LIP [huang2024CanLarge] and PromptAV [hung2023WhoWrote] with a 24B Mistral LLM as zero-shot baselines. For STAR and OSST-based methods, we select a decision boundary with 2000 examples from the training set.
@@
The results on authorship verification datasets are summarized in figs.˜4 and 5. Several key patterns emerge: (i) nOSST consistently surpasses STAR on settings that control for topical correlations (the third tasks in PAN23ST and PAN24ST), even in its zero-shot (ZS) configuration with an 8B model. In contrast, STAR is better in datasets with strong topic biases. (ii) At comparable model scales (up to 24B parameters), nOSST methods substantially outperform LLM-based prompting approaches. (iii) The RS variant of nOSST generally improves performance across multiple datasets, though a few exceptions remain. (iv) Performance exhibits a clear scaling trend: larger nOSST base models tend to achieve higher accuracy, with only minor deviations from monotonic improvement.
@@ 4.5 Multi-language performance in authorship attribution
In fig.˜8, we present the average accuracy and rank of correct candidate of the proposed OSST-based method across various languages and backbone models for authorship attribution (AA), using the PAN18 and PAN19 multi-language AA datasets. STAR is included as a baseline for comparison. As anticipated, performance declines for all models outside of English, with the OSST-based method consistently outperforming STAR by a larger margin in non-English settings. This disparity can be attributed to the limited multilingual training of STAR, whereas OSST benefits from pre-training on large-scale cross-lingual corpora. These results show OSST’s superior generalization and robust multilingual capabilities in authorship attribution tasks.
@@ 5 Conclusions
We introduced OSST-based methods, a family of unsupervised approaches to authorship attribution and verification that take advantage of the extensive causal language modeling pre-training of modern LLMs. This paradigm enables scalable authorship analysis without relying on labeled data prone to topical biases.
@@ Appendix F Do we need neutral examples?
Figure 10: Comparison between using (nOSST) and not using (nOSST-NN) a neutral version to compute log-probabilities.
Appendix G To what extent does the neutralizer model influence performance?
In this work we have used a model of 12B parameters to create neutral versions of the given texts. This choice greatly influences the amount of resources needed to run our model overall. Is this an important choice? In fig.˜11 we show a comparison between using a 4B and a 12B model in PAN22-24ST tasks. As we can see, this downgrade does affect performance slightly, but not to a great extent.
Figure 11: Comparison between using a google/gemma-3-4b-it google/gemma-3-12b-it [team2025Gemma3] model to create neutral versions of texts in PAN22ST, PAN23ST and PAN24ST.